if the sum of 5% of a and 4% of b is 2/3 of sum of 6% of a and 8%of b then find out a:b
Answer by stanbon(75887)

if the sum of 5% of a and 4% of b is 2/3 of sum of 6% of a and 8%of b then find out a:b
----
Equation:
0.05a + 0.04b = (2/3)(0.06a + 0.08b)
-----
Multiply theu by 100 to get:
5a +4b = (2/3)(6a + 8b)
----
5a + 4b = 4a + (16/3)b
----
a = (5 1/3)b - 4b
---
a = 1 1/3 b
----
a/b =4/3
===============
Cheers,
Stan H.
=====================